. Peanuts cost $0.80 a pound, cashews $1.50 a pound, anf alnuts $1.50 a pound. A store clerk needs to make 6 pound of mixed nuts using these three types so the cost is $6.90. If she wants to use 3 times as much peanuts as cashews, find how much of each type she should use.
Answer by mananth(16949)

Peanuts cost $0.80 a pound, cashews $1.50 a pound, anf alnuts $1.50 a pound. A store clerk needs to make 6 pound of mixed nuts using these three types so the cost is $6.90. If she wants to use 3 times as much peanuts as cashews, find how much of each type she should use.
Peanuts cost $0.80 a pound,..........3x pounds
cashews $1.50 a pound, ..............x pounds
and alnuts $1.50 a pound. ...........6-4x
6 pound of mixed nuts --- $6.90.
0.8*3x+1.5*x+1.5(6-4x)= 6.90
2.4x+1.5x+9-6x=6.9
3.9x-6x+9=6.9
-2.1x=-2.1
/-2.1
x=1 pound -------- cashews
Peanuuts =====> 3x====>3 pounds
alnuts=2 pounds
...